OpenStudy (anonymous):

The cost function and the demand equation for a certain product are C(x)=60x+7200, p=-2x+300 Find: a.The revenue function. b.The marginal revenue. c.The marginal cost. d.The break-even point(s). (Hint: the numbers for which R(x) = C(x).) e.The number x for which marginal revenue equals marginal cost. @cruffo

OpenStudy (dumbcow):

yeah economics revenue = price * quantity = p(x)*x marginal revenue is the derivative of revenue function... R'(x) marginal cost is derivative of cost function ...C'(x) break even point is where cost = revenue or profit is zero C(x) = R(x) ... solve for x marginal revenue = marginal cost C'(x) = R'(x) ... solve for x

OpenStudy (anonymous):

can u please write out the question a for me..

OpenStudy (dumbcow):

i did R(x) = x*p(x)

OpenStudy (anonymous):

where p(x) is -2x + 300 therefore revenue function = x(-2x + 300)?

OpenStudy (dumbcow):

yeah or -2x^2 +300x

OpenStudy (anonymous):

oh ok.so with the b i need to use the derivative formula to solve for it ..thus the d/dx

OpenStudy (dumbcow):

yes where \[\frac{d}{dx} x^{n} = nx^{n-1}\]

OpenStudy (anonymous):

yeahh kk.i have two more questions can i post one and tag u in it please

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

im stucked at the break even point question @dumbcow .can you help

OpenStudy (dumbcow):

hmm ok you should have \[60x+7200 = -2x^{2}+300x\] set quadratic equal to 0 \[2x^{2}-240x+7200 = 0\] factor out a 2 \[x^{2}-120x+3600 = 0\]

OpenStudy (anonymous):

do i use the quadratic equatiion formula here or

OpenStudy (dumbcow):

you can or try to factor it (hint its a perfect square)

OpenStudy (anonymous):

please start it for me :-)

OpenStudy (dumbcow):

well if you dont see how to factor always go to quadratic formula, that will work every time

OpenStudy (anonymous):

OH KK

OpenStudy (anonymous):

DID U SOLVE THE BREAK EVEN? IM I USING THE VERTEX H AND K OR THE the quadratic equation formula.dont know if you get me @dumbcow

OpenStudy (anonymous):

hii.i got it now...sure u out..but thanks i figured it out..